OpenStudy (anonymous):

A 1,000-kg car has 50,000 joules of kinetic energy. What is its speed?

OpenStudy (isaidavila):

The kinetic energy is 1/2mv^2 So 50000=1/2*1000v^2 100=v^2 v=10